solve this equation by substitution
x=8-4y
3x+5y=3

Answer by mananth(16949)

x+4y=8 .............1
5y=3-3x ---------2
/5=
y=0.6-0.6x
Plug the value of y in (1)
x+4(0.6 -0.6x)
x+2.4-2.4x=8
x-2.4x=8-2.4
-1.4x=5.6
/-1.4
x=-4
Plug the value of x in (1)
x+4y=8
1*-4+4y=8
-4+4y=8
4y=12
/4
y=3